OpenStudy (anonymous):

What are the solutions of x^2 - 3x +9 = 0

OpenStudy (anonymous):

Imaginary

OpenStudy (anonymous):

Thanks man.

OpenStudy (anonymous):

Sarcastic? :( I didn't technically help you.

OpenStudy (anonymous):

Well not sarcastic i figured you were at least trying to help so i appreciate that. The problem is still confusing me though? :[

OpenStudy (anonymous):

You can either work it out with the quadratic formula or completing the square, but the graph does not cross the x axis, so both solutions are imaginary.

OpenStudy (anonymous):

But i did not get two solutions in the problem? :[ im just totally confused on how to solve it :/

OpenStudy (anonymous):

OK one second:

OpenStudy (anonymous):

\[x^2 - 3x + 9 = 0\] \[\iff \left(x-\frac{3}{2}\right)^2 - \left(\frac{3}{2}\right)^2 + 9 = 0\] \[\iff \left(x-\frac{3}{2}\right)^2 = -\frac{27}{4}\] \[\implies x-\frac{3}{2} = \pm \sqrt{\frac{-27}{4}}\]

OpenStudy (anonymous):

Of course, you could equally just use the formula, but I think this way is more fun. Do you know how to carry on (mainly with the root) to get the imaginary/complex solution?

OpenStudy (anonymous):

I do not.

OpenStudy (anonymous):

Do you see how the working I posted works?

OpenStudy (anonymous):

Yeahh.....

OpenStudy (anonymous):

\[\sqrt{\frac{-27}{4}} = \sqrt{-1}\cdot\sqrt{\frac{27}{4}} = i\cdot\frac{3\sqrt{3}}{2}\]

OpenStudy (anonymous):

Again, it may be more 'standard' and easier to repeat if you use the quadratic formula instead, but it would still need you to take the root of a negative number.